# Hi Robot: Open-Ended Instruction Following with Hierarchical Vision-Language-Action Models

### v1

@@ 2 Related Work
In the latter category, a number of methods use LLMs and VLMs to reason over robot observations and commands, and break up multi-stage tasks into simpler steps that can be performed by low-level controllers. Earlier methods of this sort used language models in combination with various learned or hand-designed skills (Huang et al., 2022; Brohan et al., 2023b; Liang et al., 2023; Shah et al., 2024; Singh et al., 2023; Wang et al., 2024), but such systems have limited ability to incorporate complex context, such as image observations, into the reasoning process. More recently, multiple works have use VLMs to output parameters for pre-defined robotic skills (Huang et al., 2023; Liu et al., 2024a; Nasiriany et al., 2024; Chen et al., 2024; Liu et al., 2024b; Stone et al., 2023; Qiu et al., 2024; Zhi et al., 2024). Such methods can process more complex commands and situate them in the context of visual observations, but these approaches have shown limited physical dexterity and limited ability to incorporate real-time language interaction with humans (with some exceptions discussed below). In contrast, our system utilizes VLMs for both high-level reasoning and low-level control, with a flexible language interface between the two. These design choices, along with a new synthetic data generation scheme, allow our system to achieve both significant physical dexterity and detailed promptability that prior works lack.
@@
Many works aim to enable robotic language interaction with users, including model-based systems that parse language instructions and feedback and ground them via a symbolic representation of the scene (Swadzba et al., 2009; Matuszek et al., 2013; Namasivayam et al., 2023; Patki et al., 2019), and more recent learning-based methods that process feedback directly, typically with a hierarchical architecture (Liu et al., 2023; Xiao et al., 2024; Shi et al., 2024; Belkhale et al., 2024; Singh et al., 2024; McCallum et al., ; Driess et al., 2023; Dai et al., 2024). Our work builds on the latter class of methods, where user feedback is incorporated via a high-level policy that provides atomic commands to a learned low-level policy. Unlike OLAF (Liu et al., 2023), which uses an LLM to modify robot trajectories, our approach can incorporate situated corrections based on the robot’s observations, respond to those corrections in real time, and follow complex prompts describing dexterous manipulation tasks. While YAY Robot (Shi et al., 2024) can handle situated real-time corrections, it is limited to one prompt and to the corrections seen in the human-written data; our approach leverages VLMs and a new data generation scheme to enable diverse prompts and open-ended corrections. Finally, RACER (Dai et al., 2024) can also incorporate situated corrections, but relies on a physics simulator to construct recovery behaviors; our approach only uses real robot demonstrations without intentional perturbations or corrections and is applicable to open-ended prompts.
@@ B.2 Inference Hardware
B.3 Robot System Details

### v2

@@ 2 Related Work
In the latter category, a number of methods use LLMs and VLMs to reason over robot observations and commands, and break up multi-stage tasks into simpler steps that can be performed by low-level controllers. Earlier methods of this sort used language models in combination with various learned or hand-designed skills (Huang et al., 2022; Brohan et al., 2023b; Liang et al., 2023; Shah et al., 2024; Singh et al., 2023; Wang et al., 2024; Li et al., 2025a), but such systems have limited ability to incorporate complex context, such as image observations, into the reasoning process. More recently, multiple works have use VLMs to output parameters for pre-defined robotic skills (Huang et al., 2023; Liu et al., 2024a; Nasiriany et al., 2024; Chen et al., 2024; Liu et al., 2024b; Stone et al., 2023; Qiu et al., 2024; Zhi et al., 2024). Such methods can process more complex commands and situate them in the context of visual observations, but these approaches have shown limited physical dexterity and limited ability to incorporate real-time language interaction with humans (with some exceptions discussed below). In contrast, our system utilizes VLMs for both high-level reasoning and low-level control, with a flexible language interface between the two. These design choices, along with a new synthetic data generation scheme, allow our system to achieve both significant physical dexterity and detailed promptability that prior works lack.
@@
Many works aim to enable robotic language interaction with users, including model-based systems that parse language instructions and feedback and ground them via a symbolic representation of the scene (Swadzba et al., 2009; Matuszek et al., 2013; Namasivayam et al., 2023; Patki et al., 2019), and more recent learning-based methods that process feedback directly, typically with a hierarchical architecture (Liu et al., 2023; Xiao et al., 2024; Shi et al., 2024; Belkhale et al., 2024; Singh et al., 2024; McCallum et al., ; Driess et al., 2023; Dai et al., 2024; Hu et al., 2023; Li et al., 2025b). Our work builds on the latter class of methods, where user feedback is incorporated via a high-level policy that provides atomic commands to a learned low-level policy. Unlike OLAF (Liu et al., 2023), which uses an LLM to modify robot trajectories, our approach can incorporate situated corrections based on the robot’s observations, respond to those corrections in real time, and follow complex prompts describing dexterous manipulation tasks. While YAY Robot (Shi et al., 2024) can handle situated real-time corrections, it is limited to one prompt and to the corrections seen in the human-written data; our approach leverages VLMs and a new data generation scheme to enable diverse prompts and open-ended corrections. Finally, RACER (Dai et al., 2024) can also incorporate situated corrections, but relies on a physics simulator to construct recovery behaviors; our approach only uses real robot demonstrations without intentional perturbations or corrections and is applicable to open-ended prompts.
@@ Acknowledgments
Impact Statement
This paper presents work where the goal is to advance the fields of machine learning and robotics. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.2 Inference Hardware
B.3 Real-Time Inference Latency
We measured latency across components on consumer-grade RTX 4090.
Low-Level Policy Per-Step Inference Times
Component Time (ms)
Image encoding 14
Observation processing 32
Action prediction (x10) 27
Total (on-board) 73
Total (off-board + WiFi) 86
High-Level Policy (Single Decoding Step)
  • RTX 4090: 47 ms (prefill) + 13.2 ms (decode)
  • H100: 17.3 ms (prefill) + 5.7 ms (decode)
These measurements confirm real-time feasibility at $\sim$10 Hz control rates. With action chunking (Zhao et al., 2023), we can use it to control robots at 50 Hz.
B.4 Robot System Details
@@
Appendix C Model and Experiment Details
C.1 Model Initialization
While our method can be trained from scratch or fine-tuned from any vision-language model (VLM) backbone, in practice we use PaliGemma (Beyer et al., 2024) as the base model. PaliGemma is an open-source, 3-billion-parameter VLM that offers a good balance between performance and computational efficiency. We unfreeze the full model for fine-tuning.
C.2 Optimizer and Hyperparameters
We use the AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and no weight decay. Gradient norms are clipped to a maximum magnitude of 1. We maintain an exponential moving average (EMA) of the network weights with a decay factor of 0.999. The learning rate is warmed up over the first 1,000 steps and then held constant at $1\times 10^{-5}$. We use a batch size of 512.
C.3 Training Duration and Resources
Training the high-level policy is highly efficient, requiring approximately 2 hours on 8$\times$×H100 GPUs. The low-level policy follows a similar training pipeline, though training time may vary depending on the dataset size and the complexity of the target tasks for action prediction.
C.4 Failure Cases
We observed the following failure modes:
  1. 1.
    High-level:
    • Difficulty with instructions requiring long-context reasoning, since the current system lacks memory
  2. 2.
    Low-level:
    • Temporarily ignoring instructions: e.g., grabbing cheese when the robot is close to it despite user’s lactose intolerance (due to training bias toward proximal objects)
    • Error accumulation and out-of-distribution (OOD) recovery: e.g., dropped objects
Beyond the future directions discussed in the main text, several additional mitigations may help address observed limitations, including but not limited to:
  • Stronger instruction-following model
  • Long-context model
  • Adversarial data generation for edge cases
  • Diverse data collection including failure recovery and annotation
C.5 System Prompt for GPT-4o
In the system prompt for GPT-4o, we include a description of the task, robot setup, and common instructions to select from. Below is an example for the Table Cleaning task.
Listing 1: GPT-4o Baseline Prompt
You are an AI assistant guiding a single-arm robot to bus tables.
The robot can optionally place trash in the trash bin and utensils and dishes in the plastic box.
Every 2 seconds, you can issue one instruction from a provided list.
You will receive images from two cameras: one for a global view and one on the robot wrist for detailed views.
Interpret the users instruction into one from the provided list for the robot to execute. Adhere strictly to the users instruction. If ambiguous, reason out the best action for the robot. Only provide the exact instruction from the list without explanation.
You will select your instruction from the following list:
put food container in trash bin
pick up chopstick
drop wrapper in trash
pick up plastic plate
pick up the cup
pick up white bowl
place bowl to box
pick up spoon
place trash to trash bin
drop box in trash
place take out box to trash
move to the left
pick up container
drop plate in bin
pick up the trash
pick up plastic bowl
go higher
place spoon to box
pick up the paper container
drop fork in bin
pick up the bowl
pick up the plastic container
go lower
pick up box
move to the right
drop plastic lid into recycling bin
pick up wrapper
pick up the plate
put bowl in box
pick up the container
put the plate in the bin
pick up cup
put cup into box
throw it in the trash
pick up food container
pick up blue cup
drop the bowl into the bin
move towards me
pick up napkin
rotate counterclockwise
put the cup in the bin
throw trash away
rotate clockwise
drop plastic bowl into box
open gripper
pick up plastic cup
pick up the plate
close gripper
move away from me
go back to home position